OpenStudy (anonymous):

Which ordered pair is a solution of the equation y = –5x + 4? a-(2, 14) b-(5, –29) c-(3, –19) d-(5, –21)

OpenStudy (anonymous):

d

OpenStudy (anonymous):

y=-5(5)+4 y=-25+4 y=-21